OpenStudy (anonymous):

Write the equation of the line that passes through (–1, 5) with a slope of 2 in standard form.

OpenStudy (anonymous):

the multiple choice answers are: 2x – y = 11, –2x + y = –11, –2x – y = –7, 2x – y = –7

OpenStudy (saifoo.khan):

Sorry i copied the question incorrect. \[y = 2 x+7 \]

OpenStudy (saifoo.khan):

D looks right,